OpenStudy (anonymous):

What is the square root of 500x^3 all over the square root of 10x^-1 ??

OpenStudy (anonymous):

I got 7.0710678...(you can just round it if you want)

OpenStudy (jkristia):

do you mean to simplify like this ? \[\frac{\sqrt{500x^3}}{\sqrt{10x^{-1}}}\] this is the same as \[\sqrt{\frac{500x^3}{10x^{-1}}}\] \[\sqrt{\frac{50x^3}{x^{-1}}}\] \[\sqrt{50x^4}\] \[x^2\sqrt{50}\]